30. Januar 2002. U. Schoenwaelder; http://www.math.rwth-aachen.de/∼Ulrich.Schoenwaelder HB = Hochschulbibl. RWTH, HBZ = http://www.hbz-nrw.de/ (HBZ-CD-ROM Online), MB = Mathematikbibl., DB = Didaktikbibl. (Winter), EB = Erziehungswiss. Bibl., FH = Bibl. Fachhochschule Aachen, FL = Fernleihe, IB Nr. = Institutsbibliothek Nr., LB = HB–Lehrbuchsammlung, LS = HB–Lesesaal

Literatur über die Faktorisierung ganzer Zahlen

[1] E. Bach. Toward a theory of Pollard’s rho method. Inform. and Comput., 90:139–155, 1991.

[2] Eric Bach and Jeffrey Shallit. Factoring with cyclotomic polynomials. Math. Computation, 52(185):201–219, 1989.

[3] Eric Bach and Jeffrey Shallit. Algorithmic Number Theory, Vol I: Efficient Algorithms. Foundations of Computing. MIT Press, 1996. ISBN 0-262-02405-5. MB. HBZ.

[4] R. Bach. Number–theoretic algorithms. Ann.Rev. Comput. Sci., 4:119–172, 1990.

[5] E. Behrends. N = NP? Oder, anders gefragt: Ist Glück in der Mathematik entbehrlich? DIE ZEIT, 1999(10, 4. März):43, 1999. RSA-Verfahren.

[6] Michael N. Bleicher. The Search for Perfect Numbers. In Anatole Beck, sMichael N. Bleicher, and Donald W. Crowe, editors, Excursions into Mathematics, chapter 2, pages 79–143. A K Peters, millenium ed. 2000 edition, 2000. FL: SUB Göttingen. S. 110–124: Faktorisierung.

[7] D. Boneh. Twenty years of attacks on the RSA cryptosystem. Notices Amer. Math. Soc., 46:203–213, 1999.

[8] W. Bosma and A. K. Lenstra. An implementation of the elliptic curve integer factorization method. In W. Bosma and A. van der Poorten, editors, Computational Algebra and Number Theory, pages 119–136. Kluwer, 1995.

[9] R. Brent. Factorization of the tenth Fermat number. Math. Comp., 68:429–451, 1999.

[10] R. P. Brent. An improved Monte Carlo factorization algorithm. BIT, 20:176–184, 1980.

[11] R. P. Brent and J. M. Pollard. Factorization of the eigth Fermat number. Math. Computation, 36:627–630, 1981.

[12] D. Bressoud and S. Wagon. A Course in Computational Number Theory. Emeryville, CA: Key College Publishing, 2000. ISBN 1-930190-10-7. HBZ.

[13] David M. Bressoud. Factorization and Primality Testing. UTM. Springer–Verlag, 1989. HB: Bb1761. Lst.-B. ISBN 0-387-97040-1.

[14] J. Brillhart, D. H. Lehmer, J. L. Selfridge, B. Tuckerman, and jr. S. S. Wagstaff. Factorizations of b^{n} ± 1, b = 2, 3, 5, 6, 7, 10, 11, 12 up to high powers. Contemporary Mathematics 22. AMS, 1983. Einführung, Übersicht, großes Literaturverzeichnis.

[15] Stefania Cavallar et al. Factorization of a 512-bit RSA modulus. In B. Preneel, editor, Advances in Cryptology – EURO-CRYPT 2000 (Intern. Conf. .., Bruges, Belgium, May 2000), number 1807 in LNCS, page 1 ff. Springer–Verlag, 2000. http://link.springer.de/ see Books/Book Series.

[16] H. Cohen. A Course in Computational Algebraic Number Theory. GTM 138. Springer–Verlag, 1993. Ch. 8: Factoring in the dark ages; Ch. 10: Modern factoring methods.

[17] Richard Crandall and Carl Pomerance. Prime Numbers: A Computational Perspective. New York: Springer, 2001. ISBN 0-387-94777-9. MB: 19315. 1 Primes! 1.1 Problems and progress; 1.2 Celebrated conjectures and curiosities; 1.3 Primes of special form; 1.4 Analytic number theory. 2 Number-theoretical tools 2.1 Modular arithmetic; 2.2 Polynomial arithmetic; 2.3 Squares and roots. 3 Recognizing primes and composites 3.1 Trial division; 3.2 Sieving; 3.3 Pseudoprimes; 3.4 Probable primes and witnesses; 3.5 Lucas pseudoprimes; 3.6 Counting primes. 4 Primality proving 4.1 The n − 1 test; 4.2 The n + 1 test; 4.3 The finite field primality test; 4.4 Gauss and Jacobi sums. 5 Exponential factoring algorithms 5.1 Squares; 5.2 Monte Carlo methods; 5.3 Baby-steps, giant-steps; 5.4 Pollard p − 1 method; 5.5 Binary quadratic forms. 6 Subexponential factoring algorithms 6.1 The quadratic sieve factoring algorithm; 6.2 Number field sieve; 6.3 Rigorous factoring; 6.4 Index-calculus method for discrete logarithms. 7 Elliptic curve arithmetic. 8 The ubiquity of prime numbers 8.1 Crytography .. 9 Fast algorithms for large-integer arithmetic. S. 501–525: References.

[18] H. Davenport. The Higher Arithmetic. An Introduction to the Theory of Numbers. Hutchinson & Co. Ltd.; Cambridge Univ. Press, 1952, ^{6}1992. MB: 16 723. Inhalt: I Factorization and primes. II Congruences. III Quadratic residues. IV Continued fractions. V Sums of squares. VI Quadratic forms. VII Some Diophantine equations. VIII Computers and the theory of numbers (4. Pollard’s factoring method. 5. Factoring large numbers. 7. The RSA cryptographic method).

[19] K. Devlin. Sternstunden der modernen Mathematik. Berühmte Probleme und neue Lösungen. Birkhäuser, 1990. MB: 15864. Die Originalausgabeerschien 1988 unter dem Titel ”Mathematics: The New Golden Age“ bei Penguin Books Ltd. §1: Primzahlen, Faktorzerlegung und Geheimcodes.

[20] J. D. Dixon. Factorization and primality tests. Amer. Math. Monthly, 91(6):333–352, 1984. MB.

[21] jr. H. W. Lenstra. Factoring integers with elliptic curves. Annals of Math., 126:649–673, 1987.

[22] H. Kleiner. Das RSA–Verfahren. Praxis der Mathematik, 26(5):133–140, 1984. HB: Z1757-26. MB: Z 101.

[23] N. Koblitz. A Course in Number Theory and Cryptography. GTM 114. Springer–Verlag, 1987. HB: 114 Bb 1244. Ch. I: Some topics in elementary number theory; §1: Time estimates for doing arithmetic, §2: Divisibility and the Euclidean algorithm, §3: Congruences (Fermat’s Little Theorem, Chinese Remainder Theorem), §4: Some applications to factoring. Ch. II: Finite fields and quadratic residues. Ch. III: Cryptography. Ch. IV: Public Key. Ch. V: Primality and Factoring. Ch. VI: Elliptic Curves (cryptosystems, factorization).

[24] D. H. Lehmer. Factorization then and now. In D. V. Chudnovsky and R. D. Jenks, editors, Computers in Mathematics, LNPAM 125. Dekker, 1990.

[25] A. K. Lenstra and H. W. Lenstra, jr. Algorithms in number theory. In J. van Leeuwen, editor, Handbook of Theoretical Computer Science, Vol. A, pages 673–715. Elsevier, Amsterdam, 1990.

[26] A. K. Lenstra, H. W. Lenstra, jr., M. S. Manasse, and S. M. Pollard. The factorization of the ninth Fermat number. Preprint, 1991.

[27] A. K. Lenstra and H. W. Lenstra, editors. The Development of the Number Field Sieve, LNM 1554. Springer–Verlag, 1993. Six research papers on the number field sieve for factoring large integers. A variant applies in general.

[28] H. W. Lenstra and C. Pomerance. A rigorous time bound for factoring integers. Journal of the AMS, 5:483–516, 1992. MR 92m: 11145.

[29] P. L. Montgomery. Speeding the Pollard and elliptic curve methods of factorization. Math. Computation, 48(177):243–264, 1987.

[30] M. A. Morrison and J. Brillhart. A method of factoring and the factorization of F_{7}. Math. Computation, 29(129):183–205, 1975.

[31] Cornelia Niederdrenk-Felgner. Das RSA-Verschlüsselungssystem – Ein Thema für die Schule? Beiträge zum Mathemsatikunterricht, 1989:274–277, 1989. HB: Bb1256-1989.

[32] J. M. Pollard. Theorems on factorization and primality testing. Proc. Camb. Philos. Soc., 76:521–528, 1974.

[33] C. Pomerance. Analysis and comparison of some integer factoring algorithms. In H. Lenstra and R. Tijdeman, editors, Computational Methods in Number Theory, pages 89–141. 1982.

[34] C. Pomerance. Implementation of the continued fraction integer factoring algorithm. Congress. Numer., 37:99–118, 1983.

[35] C. Pomerance. The quadratic sieve factoring algorithm. In T. Beth, N. Cot, and I. Ingemanson, editors, Advances in Cryptology, LNCS 209, pages 169–182. Springer–Verlag, 1985.

[36] C. Pomerance. Factoring. In C. Pomerance, editor, Cryptology and Computational Number Theory, Proc. of Symposia in Appl. Math. 42, pages 27–47. AMS, 1990.

[37] H. I. Riesel. Prime Numbers and Computer Methods of Factorization. PM 126. Birkhäuser, 1985, ^{2}1994. MB: 12884.

[38] R. Rumely. Recent advances in primality testing. Notices AMS, 30(5):475–477, 1985. Übersichtsartikel zu Primzahltests und Faktorisierung.

[39] S. S. Wagstaff, jr. and J. W. Smith. Methods of factoring large integers. In D. V. Chudnovsky, G. V. Chudnovsky, H. Cohn, and M. B. Nathanson, editors, Number Theory (New York, 1984-1985), LNM 1240, pages 281–303. Springer–Verlag, 1987. Übersichtsartikel.

[40] R. D. Silverman. The multiple polynomial quadratic sieve. Math. Computation, 48(177):329–339, 1987.

[41] M. C. Wunderlich. Implementing the continued fraction factoring algorithm on parallel machines. Math. of Computation, 44(169):251–260, 1985.